est to determine if there is a significant difference in the average cost of textbooks for 3 students whose textbook costs are noted below. Alexia 160 100 220 120 150 Demi 180 120 180 110 150 200 Shikha 150 180 150 160 120 What is the value of the Error MS = ______ (round to the third value to the right of the decimal)
Answer #1

Here I used Excel to compute the statistical results.

the given problem is one way anova

SS(Total) = SS(Between) + SS(Error)

SS(E)=SS(TO)−SS(T)

m ni

= 16813.33

MSE= 16813.33/13 = 1293.333

16813.33

SUMMARY

Groups

Count

Sum

Average

Variance

Alexia

5

750

150

2100

Demi

6

940

156.6667

1306.667

Shikha

5

760

152

470

ANOVA

Source of Variation

SS

df

MS

F

P-value

F crit

Between Groups

130.4167

2

65.20833

0.050419

0.951016

3.805565

Within Groups

16813.33

13

1293.333

Total

16943.75

15